Tin that was 48% pure was melted together with tin that was 44% pure to make 20 kilograms of an alloy that was 45% pure. How many kilograms of each were used? How do you set it up?
Answer by Paul(988)

0.48x+0.44(20-x)=0.45(20)
0.48+8.8-0.44x=9
0.04x=0.2
x=5

20-5=15

Hence, for 48% about 5 kg was used and for 44% about 15 kg was used.
Paul.